OpenStudy (anonymous):

can u guys help me with this augmented matrice that have no solutions, many solutions and unique solutions?

OpenStudy (anonymous):

question number 2..

OpenStudy (turingtest):

If I understand the first one correctly we want to solve for the number of small, medium, and large units as a function of the sub-units A, B, and C. It looks like we will turn this matrix sideways then: small: 2A+B+C=6 medium:4A+3B+5C=12 large: 9A+5B+8C=15 so we make a matrix from the coefficientsnow perform Cramer's rule on this system to find A, B, and C